OpenStudy (anonymous):

Determine the minimum of the following function. Y = x2 – 14x + 6

OpenStudy (anonymous):

7

OpenStudy (anonymous):

calculus way or algebra? lol

OpenStudy (akshay_budhkar):

you can use the calculus way differentiate it equal to 0 get the values double diffrentiate to get minimum and max. value at what value of x and substitute it in the original equation

OpenStudy (akshay_budhkar):

you get it? or did i sound greek?

OpenStudy (anonymous):

vertex = 14/2 = 7 value of function at x=7 = 49 - 14(7) + 6 = -43 = minimum value